OpenStudy (anikate):

the ratio of 4 less than a number to 26 more than that number is 1 to 3. What is the number?

OpenStudy (anonymous):

represent the number as x, write the ratio as a fraction with an expression in the numerator and denominator and then set it equal to 1/3. cross multiply and solve for x

OpenStudy (anonymous):

e.g. if we call a number x, what is an expression for 4 less than that number x, and what is another expression for 26 more than that number x

OpenStudy (anikate):

i dont get it?

OpenStudy (anonymous):

ok, so if I say I have 4 less than 25 dollars, I have 25-4 dollars=21, right? So 4 less than x would be...

OpenStudy (anikate):

4-x

OpenStudy (anonymous):

right, so 26 more than x is x +26

OpenStudy (anonymous):

and the ratio of two expressions can be written as a fraction so 4-x/x+26

OpenStudy (anikate):

that's the answer

OpenStudy (anonymous):

well almost, you need to set it equal to the ration they give you and then solve for x

OpenStudy (anonymous):

ratio*

OpenStudy (anikate):

so it would be?

OpenStudy (anonymous):

4-x/x+26=1/3 Cross multiply to get 3(4-x)=x+26. 12-3x=x+26, 2x=38, x=19

OpenStudy (anikate):

thanks man

OpenStudy (anonymous):

np